# Busemann Functions in the Wasserstein Space: Existence, Closed-Forms, and Applications to Slicing

### v1

@@ B.1 Background on the Wasserstein over Bures-Wasserstein Space
Let $\mathbb{P}\in\mathcal{P}_{2}\big(\mathrm{BW}(\mathbb{R}^{d})\big)$, $\theta\in S^{d-1}$ and $\varphi^{\theta}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathcal{P}_{2}(\mathbb{R})$ defined as $\varphi^{\theta}(\mu)=P^{\theta}_{\#}\mu$. For any $\mu=\mathcal{N}(m_{\mu},\Sigma_{\mu})\sim\mathbb{P}$, $\varphi^{\theta}(\mu)=\mathcal{N}(\langle m_{\mu},\theta\rangle,\theta^{T}\Sigma_{\mu}\theta)$. Thus, $\mathbb{P}^{\theta}=\varphi^{\theta}_{\#}\mathbb{P}\in\mathcal{P}_{2}\big(\mathrm{BW}(\mathbb{R})\big)$. Moreover, the space of 1D Gaussian can be identified as a product space over the means and standard deviations $\mathbb{R}\times\mathbb{R}_{+}^{*}$. For $\mu_{\theta}=\mathcal{N}(m_{\theta},\Sigma_{\theta})\in\mathrm{BW}(\mathbb{R})$, let $\Xi:\mathrm{BW}(\mathbb{R})\to\mathbb{R}^{2}$ such that $\Xi(\mu_{\theta})=(m_{\theta},\sigma_{\theta})$. Then $\Xi_{\#}\mathbb{P}^{\theta}\in\mathcal{P}_{2}(\mathbb{R}^{2})$. Nguyen and Mueller [2024], Piening and Beinert [2025a] proposed to define the Doubly Mixture Sliced-Wasserstein distance (DMSW) between $\mathbb{P},\mathbb{Q}\in\mathcal{P}_{2}\big(\mathrm{BW}(\mathbb{R}^{d})\big)$ as
@@ D.1.5 Proof of Corollary 3
Let $\mu_{0}=\mathcal{N}(m_{0},\Sigma_{0})$ and $\mu_{1}=\mathcal{N}(m_{1},\Sigma_{1})$ with $m_{0},m_{1}\in\mathbb{R}^{d}$ and $\Sigma_{0},\Sigma_{1}$ symmetric positive definite matrices. The Monge map between $\mu_{0}$ and $\mu_{1}$ is [Peyré et al., 2019, Remark 2.31]
@@ F.1 Clustering Detection
We perform the same experiment as in [Piening and Beinert, 2025a, Section 4.3]. We want to fit a Gaussian mixture on a dataset. A common technique is to use an EM. However, the number of clusters in the mixture needs to be specified and is not necessarily known beforehand. For $K$ components, denote $\mathbb{P}_{K}\in\mathcal{P}_{2}\big(\mathrm{BW}(\mathbb{R}^{d})\big)$ the corresponding mixture. Piening and Beinert [2025a] propose to compute the distances between $\mathbb{P}_{k}$ and $\mathbb{P}_{k+1}$ for $k\geq 1$, and to find the smallest $k\in\mathbb{N}^{*}$ such that adding more components does not change the fitted model. To do so, they increase $k$ until the distances vanish.
We report on Figure 11 the results using the same setting as [Piening and Beinert, 2025a], i.e. we have 1500 samples of a mixture in dimension $d=2$ with 4 clusters. On Figure 12, we add an experiment in dimensions $d\in\{3,5,10,20\}$ with $K\in\{4,5,6\}$ components, with means distributed uniformly on a sphere with radius $r=100$ and random covariance matrices. In every cases, we observe that the results between DSMW, $\mathrm{BGMSW}$ and $\mathrm{B1DGMSW}$ are very close.
For DMSW, we use the code of [Piening and Beinert, 2025a], available at https://github.com/MoePien/sliced_OT_for_GMMs and for $\mathrm{W}_{\mathrm{BW}}$, we use the implementation of POT [Flamary et al., 2021].

### v2

@@ B.1 Background on the Wasserstein over Bures-Wasserstein Space
Let $\mathbb{P}\in\mathcal{P}_{2}\big(\mathrm{BW}(\mathbb{R}^{d})\big)$, $\theta\in S^{d-1}$ and $\varphi^{\theta}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathcal{P}_{2}(\mathbb{R})$ defined as $\varphi^{\theta}(\mu)=P^{\theta}_{\#}\mu$. For any $\mu=\mathcal{N}(m_{\mu},\Sigma_{\mu})\sim\mathbb{P}$, $\varphi^{\theta}(\mu)=\mathcal{N}(\langle m_{\mu},\theta\rangle,\theta^{T}\Sigma_{\mu}\theta)$. Thus, $\mathbb{P}^{\theta}=\varphi^{\theta}_{\#}\mathbb{P}\in\mathcal{P}_{2}\big(\mathrm{BW}(\mathbb{R})\big)$. Moreover, the space of 1D Gaussian can be identified as a product space over the means and standard deviations $\mathbb{R}\times\mathbb{R}_{+}^{*}$. For $\mu_{\theta}=\mathcal{N}(m_{\theta},\Sigma_{\theta})\in\mathrm{BW}(\mathbb{R})$, let $\Xi:\mathrm{BW}(\mathbb{R})\to\mathbb{R}^{2}$ such that $\Xi(\mu_{\theta})=(m_{\theta},\sigma_{\theta})$. Then $\Xi_{\#}\mathbb{P}^{\theta}\in\mathcal{P}_{2}(\mathbb{R}^{2})$. Piening and Beinert [2025], Nguyen and Mueller [2026] proposed to define the Doubly Mixture Sliced-Wasserstein distance (DMSW) between $\mathbb{P},\mathbb{Q}\in\mathcal{P}_{2}\big(\mathrm{BW}(\mathbb{R}^{d})\big)$ as
@@ D.1.5 Proof of Corollary 3
Let $\mu_{0}=\mathcal{N}(m_{0},\Sigma_{0})$ and $\mu_{1}=\mathcal{N}(m_{1},\Sigma_{1})$ with $m_{0},m_{1}\in\mathbb{R}^{d}$ and $\Sigma_{0},\Sigma_{1}$ symmetric positive definite matrices. The Monge map between $\mu_{0}$ and $\mu_{1}$ is [Peyré and Cuturi, 2019, Remark 2.31]
@@ F.1 Clustering Detection
We perform the same experiment as in [Piening and Beinert, 2025, Section 4.3]. We want to fit a Gaussian mixture on a dataset. A common technique is to use an EM. However, the number of clusters in the mixture needs to be specified and is not necessarily known beforehand. For $K$ components, denote $\mathbb{P}_{K}\in\mathcal{P}_{2}\big(\mathrm{BW}(\mathbb{R}^{d})\big)$ the corresponding mixture. Piening and Beinert [2025] propose to compute the distances between $\mathbb{P}_{k}$ and $\mathbb{P}_{k+1}$ for $k\geq 1$, and to find the smallest $k\in\mathbb{N}^{*}$ such that adding more components does not change the fitted model. To do so, they increase $k$ until the distances vanish.
We report on Figure 11 the results using the same setting as [Piening and Beinert, 2025], i.e. we have 1500 samples of a mixture in dimension $d=2$ with 4 clusters. On Figure 12, we add an experiment in dimensions $d\in\{3,5,10,20\}$ with $K\in\{4,5,6\}$ components, with means distributed uniformly on a sphere with radius $r=100$ and random covariance matrices. In every cases, we observe that the results between DSMW, $\mathrm{BGMSW}$ and $\mathrm{B1DGMSW}$ are very close.
For DMSW, we use the code of [Piening and Beinert, 2025], available at https://github.com/MoePien/sliced_OT_for_GMMs and for $\mathrm{W}_{\mathrm{BW}}$, we use the implementation of POT [Flamary et al., 2021].